# Steady state of fast-oscillating neutrinos in an inhomogeneous medium

### v1

@@ I Introduction
This paper is organized as follows. In Sec. II, we introduce the neutrino equations of motion and investigate the role of the advective term in such equations, as well as its interplay with fast flavor instabilities. Our system setup is outlined in Sec. III. Then, we present our findings on the impact of advection on flavor evolution in the linear and non-linear regimes in Sec. IV. We forecast the steady-state flavor configuration and explore the dependence of the latter on the features of the flavor-dependent angular distributions in Sec. V. We discuss our findings in Sec. VI, before concluding in Sec. VII.
@@ II.2 Neutrino streaming
where $\langle\mathbf{D}_{l}\rangle\equiv\mathbf{\hat{D}}_{n}(k_{m}=0)$ denotes spatially averaged quantities across the box. The second term on the right-hand side of Eq. II.2 couples all Fourier modes and changes both direction and module of each $\mathbf{\hat{D}}_{v}(k_{m})$. Hence, from here, we see that coherent interactions can stimulate the growth of short-length inhomogeneities.
@@ III System setup
Table 1: Parameters of our benchmark angular distributions; see Eq. 8 and the right panel of Fig. 1.
@@ IV.1 Linear regime
Figure 1: Left: Growth rate of flavor instabilities as a function of their effective wavenumber, $K=k+\langle D^{z}_{1}\rangle$, where $\langle D_{1}^{z}\rangle$ is the space-averaged ELN-XLN flux, for the angular distributions C1–C4 defined by Eqs. (8) and Table 1. The vertical dotted (dashed) lines signal the location of the homogeneous (fastest-growing) mode. The dots on each $\omega_{I}(k)$ curve mark the Fourier modes $k_{m}=2\pi m/L$ for $0\leq m\leq 30$; this highlights that periodic boundary conditions impose discrete spectra on the solutions of the equations of motion. Right: Maximum growth rate of the flavor instability for the single-crossed angular distributions defined by Eqs. (8), in the plane spanned by the space-averaged ELN-XLN ($\langle D_{0}^{z}\rangle\propto\alpha$, cf. Eq. 9) and the width of the $\bar{\nu}_{e}$ distribution ($\sigma$). For all cases in this parameter space, the fastest-growing mode is inhomogeneous ($k\neq 0$).
The left panel of Fig. 1 shows the growth rate of the flavor instability as a function of the effective wavenumber of the perturbation, $K\equiv k+\langle D_{1}^{z}\rangle$, for the ELN-XLN configurations C1–C4 summarized in Table 1. The behavior of these four (anti)neutrino configurations in the linear regime is dominated by inhomogeneous modes: $\omega_{I}(K)$ has a maximum at $K_{\mathrm{max}}\neq\langle D_{1}^{z}\rangle$, which is an order of magnitude higher than the homogeneous growth rate (cf. dotted vs. dashed vertical lines in Fig. 1). Deeper angular crossings (smaller $\sigma$) correlate with higher $\omega_{I}$ across all wavenumbers and a milder dependence on $K$. A smaller ELN-XLN (higher $\alpha$) leads to a higher $\omega_{I}$ when $K\neq\langle D_{1}^{z}\rangle$. This trend is generalized in the right panel of Fig. 1, which shows the maximum growth rate for each flavor configuration in our suite of angular distributions (cf. Sec. III and Eq. 8).
@@ IV.2 Non-linear regime
Early in the non-linear phase, the evolution of the ensemble mildly depends on the shape of the initial perturbation (results not shown here). Equation (10) describes a Gaussian centered around $r=L/2$. Flavor instabilities first manifest in the middle of the box, in the proximity of $L/2$, and then spread due to advection. In the vicinity of the initial perturbation, non-linear flavor conversion begins earlier than in other regions of the box. Here, the Fourier spectrum of the ensemble is still relatively narrow and dominated by the homogeneous mode, so $D_{0}^{z}$ oscillates almost periodically shortly after the onset of flavor conversion, as shown in the right panel of Fig. 2. This pendulum-like behavior is destroyed as the coupling between length scales in Eq. (II.2) continues to broaden the spectrum. For spatial locations farther away from the peak of the initial perturbation, which are reached by the time inhomogeneities have cascaded down to smaller scales, the non-linear regime of flavor conversion begins later and no periodic features are observed in $D_{0}^{z}$ (cf. dark vs. gray lines in the right panel of Fig. 2).
@@ V Forecast of the steady-state flavor configuration
where $I_{\pm}=\int_{-1}^{+1}\mathrm{d}v\,\Theta(\pm\langle D_{v}^{z}(0)\rangle)\langle D_{v}^{z}(0)\rangle$, and $\Theta(x)$ is the Heaviside theta function. The left panel of Fig. 4 shows $\langle P_{ee}(t_{f})\rangle$, obtained by averaging Eq. 13 over the simulation box. The white dashed contour highlights the validity range of Eq. 14. In particular, we plot the relative error $\Delta_{\rm{steady}}=|\langle P_{ee}(t_{f})\rangle-\langle P_{ee}(t_{f})\rangle_{\rm{steady}}|/\langle P_{ee}(t_{f})\rangle$, and highlight the region with $\Delta_{\rm{steady}}\lesssim 8\%$. We can see that Eq. 14 performs extremely well and allows us to predict the final flavor configuration without solving the neutrino kinetic equations.
@@
Equation 14 generalizes two possible flavor outcomes, depending on $\langle D_{0}^{z}\rangle$.
@@
This criterion corresponds to the dashed cyan contour in the left panel of Fig. 4 (the dotted contour refers to a relative error $\lesssim 8\%$) and holds for our benchmark cases C1 and C2; see also the right panel of Fig. 4. We find that Eq. 16 only holds for $\langle D_{0}^{z}\rangle>0.25$ for our ensemble of neutrino angular distributions. Therefore, equipartition on one side of the angular crossing is not a generic outcome of fast flavor conversion, unlike suggested in Refs. [33, 76, 56, 57, 77].
@@ VI Discussion
By comparing cases C1–C4 in Fig. 4, we can see that the steady-state flavor configuration depends on the ELN-XLN number ($\langle D_{0}^{z}\rangle$) and the slope of the ELN-XLN in the proximity of $v_{c}$. Approximating the shape of the ELN-XLN distribution may require a large number of multipoles $\mathbf{D}_{l}$ even in the absence of flavor conversion [27, 78]; collective instabilities in inhomogeneous media should be expected to generate non-trivial contributions at higher orders, according to Eq. (5b). In this sense, attempts to find the steady state of fast flavor conversion by truncating the multipole expansion of the density matrices at an arbitrarily low $l$ cannot be predictive. For example, Ref. [79] considers $\mathbf{D}_{l}=0$ for $l>3$; we have tested that their prescription does not qualitatively reproduce the survival probabilities shown in Fig. 4 (results not shown here). This is due to the fact that our initial ELN-XLN distributions include non-trivial higher-order multipoles ($l>3$), which grow significantly over time in all our selected examples (cf. Fig. 3). This finding is in agreement with the conclusions reached in Refs. [72, 14] for a homogeneous system. Moreover, approximate solutions of Eq. (5b) were derived in Ref. [79] under the assumption that the polarization vectors are parallel to the $z$-axis at late times, which is not generally true (cf. Fig. 3, bottom panels).
@@
i.e., that the convolution of inhomogeneous modes on the right-hand side of Eq. (II.2) is negligible. We have generalized the approach introduced in Ref. [80], which considers two neutrino beams colliding head-on ($v=\pm 1$), to the case of a continuous angular distribution. We find that Eq. (17) holds at all times for initial configurations with shallow crossings and large neutrino-antineutrino asymmetries. However, when the ELN-XLN distribution is more forward peaked, Eq. (17) is only accurate at very early and very late times (results now shown here). This finding can be explained by Fig. 3, where it is clear that perturbations with large $k$ tend to subside at late times due to smearing effects induced by advection, as also numerically found in Refs. [65, 66].
@@ VII Conclusions
We offer a simple empirical formula to forecast the steady-state flavor configuration in the aftermath of flavor conversion in an inhomogeneous system. In particular, for initial ELN-XLN configurations with large neutrino-antineutrino asymmetry ($\langle D_{0}^{z}\rangle\gtrsim 0.25$), we find that the angular crossing is erased by flavor conversion. However, for small neutrino-antineutrino asymmetries ($\langle D_{0}^{z}\rangle\lesssim 0.20$) and more pronounced forward-peaked distributions, new crossings can be created as the “flipped” region of the angular distribution changes sign. For small neutrino-antineutrino asymmetry ($\langle D_{0}^{z}\rangle\lesssim 0.25$), flavor evolution is responsible for driving the system towards flavor equipartition. Our results suggest that the final flavor configuration strongly depends on $\langle D_{0}^{z}\rangle$. Thus, it cannot be predicted relying on arbitrary truncations of the multipole expansion of $D_{v}^{z}$. These findings generalize earlier approaches proposed in the literature, highlighting their limitations.
@@ Acknowledgements.
This project has received support from the European Union (ERC, ANET, Project No. 101087058) and the German Research Foundation (DFG) through the Collaborative Research Center “Neutrinos and Dark Matter in Astro- and Particle Physics (NDM),” Grant No. SFB-1258-283604770. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The Tycho supercomputer hosted at the SCIENCE HPC Center at the University of Copenhagen was used to support our numerical simulations.

### v2

@@ I Introduction
This paper is organized as follows. In Sec. II, we introduce the neutrino equations of motion and investigate the role of the advective term in such equations, as well as its interplay with fast flavor instabilities. Our system setup is outlined in Sec. III. Then, we present our findings on the impact of advection on flavor evolution in the linear and non-linear regimes in Sec. IV. We forecast the steady-state flavor configuration and explore the dependence of the latter on the features of the flavor-dependent angular distributions in Sec. V. We discuss our findings in Sec. VI, before concluding in Sec. VII. Finally, we explore the impact of the initial perturbation on the steady-state flavor configuration in Appendix A.
@@ II.2 Neutrino streaming
where $\langle\mathbf{D}_{l}\rangle\equiv\mathbf{\hat{D}}_{n}(k_{m}=0)$ denotes spatially averaged quantities across the box. The second term on the right-hand side of Eq. (II.2) couples all Fourier modes and changes both direction and module of each $\mathbf{\hat{D}}_{v}(k_{m})$. Hence, from here, we see that coherent interactions can stimulate the growth of short-length inhomogeneities.
@@ III System setup
Table 1: Parameters of our benchmark angular distributions; see Eq. (8) and the right panel of Fig. 1.
@@ IV.1 Linear regime
Figure 1: Left: Growth rate of flavor instabilities as a function of their effective wavenumber, $K=k+\langle D^{z}_{1}\rangle$, where $\langle D_{1}^{z}\rangle$ is the space-averaged ELN-XLN flux, for the angular distributions C1–C4 defined by Eqs. (8) and Table 1. The vertical dotted (dashed) lines signal the location of the homogeneous (fastest-growing) mode. The dots on each $\omega_{I}(k)$ curve mark the Fourier modes $k_{m}=2\pi m/L$ for $0\leq m\leq 30$; this highlights that periodic boundary conditions impose discrete spectra on the solutions of the equations of motion. Right: Maximum growth rate of the flavor instability for the single-crossed angular distributions defined by Eqs. (8), in the plane spanned by the space-averaged ELN-XLN ($\langle D_{0}^{z}\rangle\propto\alpha$, cf. Eq. (9)) and the width of the $\bar{\nu}_{e}$ distribution ($\sigma$). For all cases in this parameter space, the fastest-growing mode is inhomogeneous ($k\neq 0$).
The left panel of Fig. 1 shows the growth rate of the flavor instability as a function of the effective wavenumber of the perturbation, $K\equiv k+\langle D_{1}^{z}\rangle$, for the ELN-XLN configurations C1–C4 summarized in Table 1. The behavior of these four (anti)neutrino configurations in the linear regime is dominated by inhomogeneous modes: $\omega_{I}(K)$ has a maximum at $K_{\mathrm{max}}\neq\langle D_{1}^{z}\rangle$, which is an order of magnitude higher than the homogeneous growth rate (cf. dotted vs. dashed vertical lines in Fig. 1). Deeper angular crossings (smaller $\sigma$) correlate with higher $\omega_{I}$ across all wavenumbers and a milder dependence on $K$. A smaller ELN-XLN (higher $\alpha$) leads to a higher $\omega_{I}$ when $K\neq\langle D_{1}^{z}\rangle$. This trend is generalized in the right panel of Fig. 1, which shows the maximum growth rate for each flavor configuration in our suite of angular distributions (cf. Sec. III and Eq. (8)).
@@ IV.2 Non-linear regime
Early in the non-linear phase, the evolution of the ensemble mildly depends on the shape of the initial perturbation, however the steady-state flavor configuration is negligibly affected by the choice of the perturbation, as discussed in Appendix A. Equation (10) describes a Gaussian centered around $r=L/2$. Flavor instabilities first manifest in the middle of the box, in the proximity of $L/2$, and then spread due to advection. In the vicinity of the initial perturbation, non-linear flavor conversion begins earlier than in other regions of the box. Here, the Fourier spectrum of the ensemble is still relatively narrow and dominated by the homogeneous mode, so $D_{0}^{z}$ oscillates almost periodically shortly after the onset of flavor conversion, as shown in the right panel of Fig. 2. This pendulum-like behavior is destroyed as the coupling between length scales in Eq. (II.2) continues to broaden the spectrum. For spatial locations farther away from the peak of the initial perturbation, which are reached by the time inhomogeneities have cascaded down to smaller scales, the non-linear regime of flavor conversion begins later and no periodic features are observed in $D_{0}^{z}$ (cf. dark vs. gray lines in the right panel of Fig. 2).
@@ V Forecast of the steady-state flavor configuration
where $I_{\pm}=\int_{-1}^{+1}\mathrm{d}v\,\Theta(\pm\langle D_{v}^{z}(0)\rangle)\langle D_{v}^{z}(0)\rangle$, and $\Theta(x)$ is the Heaviside theta function.
The left panel of Fig. 4 shows $\langle P_{ee}(t_{f})\rangle$, obtained by averaging Eq. (13) over the simulation box. The white dashed contour highlights the validity range of Eq. (14). In particular, we plot the relative error $\Delta_{\rm{steady}}=|\langle P_{ee}(t_{f})\rangle-\langle P_{ee}(t_{f})\rangle_{\rm{steady}}|/\langle P_{ee}(t_{f})\rangle$ and highlight the region with $\Delta_{\rm{steady}}\lesssim 8\%$. We can see that Eq. (14) performs extremely well and allows us to predict the final flavor configuration without solving the neutrino kinetic equations.
@@
We compare Eq. (14) with two alternative prescriptions, which have been considered in the literature to predict the flavor outcome. We find that the agreement of such prescriptions with our numerical calculations strongly depends on $\langle D_{0}^{z}\rangle$.
@@
This criterion corresponds to the dashed cyan contour in the left panel of Fig. 4 (the dotted contour refers to a relative error $\lesssim 8\%$) and holds for our benchmark cases C1 and C2; see also the right panel of Fig. 4. We find that Eq. (16) only holds for $\langle D_{0}^{z}\rangle>0.25$ for our ensemble of neutrino angular distributions. Therefore, equipartition on one side of the angular crossing is not a generic outcome of fast flavor conversion, unlike suggested in Refs. [33, 76, 56, 57, 77].
@@ VI Discussion
By comparing cases C1–C4 in Fig. 4, we can see that the steady-state flavor configuration depends on the ELN-XLN number ($\langle D_{0}^{z}\rangle$) and the slope of the ELN-XLN in the proximity of $v_{c}$. In particular, Eq. (14) accounts for the fact that the areas of the angle-integrated ELN-XLN distribution for $v<v_{c}$ and $v>v_{c}$ are responsible for determining the steady-state configuration based on conservation laws. Approximating the shape of the ELN-XLN distribution may require a large number of multipoles $\mathbf{D}_{l}$ even in the absence of flavor conversion [27, 78]; collective instabilities in inhomogeneous media should be expected to generate non-trivial contributions at higher orders, according to Eq. (5b). In this sense, attempts to find the steady state of fast flavor conversion by truncating the multipole expansion of the density matrices at an arbitrarily low $l$ cannot be predictive. For example, Ref. [79] considers $\mathbf{D}_{l}=0$ for $l>3$; we have tested that their prescription does not qualitatively reproduce the survival probabilities shown in Fig. 4 (results not shown here). This is due to the fact that our initial ELN-XLN distributions include non-trivial higher-order multipoles ($l>3$), which grow significantly over time in all our selected examples (cf. Fig. 3). This finding is in agreement with the conclusions reached in Refs. [72, 14] for a homogeneous system. Moreover, approximate solutions of Eq. (5b) were derived in Ref. [79] under the assumption that the polarization vectors are parallel to the $z$-axis at late times, which is not generally true (cf. Fig. 3, bottom panels).
@@
i.e., that the convolution of inhomogeneous modes on the right-hand side of Eq. (II.2) is negligible. We have generalized the approach introduced in Ref. [80], which considers two neutrino beams colliding head-on ($v=\pm 1$), to the case of a continuous angular distribution. We find that Eq. (17) holds at all times for initial configurations with shallow crossings and large neutrino-antineutrino asymmetries. However, when the ELN-XLN distribution is more forward peaked, Eq. (17) is only accurate at very early and very late times (results not shown here). This finding can be explained by Fig. 3, where it is clear that perturbations with large $k$ tend to subside at late times due to smearing effects induced by advection. This results reminds us of the damping of small-scale structures numerically found in Refs. [65, 66]; however, we note that Refs. [65, 66] adopted variable neutrino densities across the simulation shell in the presence of advection, allowing for the spreading of large-scale inhomogeneities that are not present in the system setup investigated here.
@@ VII Conclusions
We offer a simple empirical formula to forecast the steady-state flavor configuration in the aftermath of flavor conversion in an inhomogeneous system. We aim to predict the angle-integrated survival probability $\langle P_{ee}(t_{f})\rangle$ averaged over our simulation box, since this is the quantity expected to be of relevance for assessing the feedback of neutrino conversion on the source physics. However, we stress that our recipe for $\langle P_{ee}(t_{f})\rangle$ intrinsically depends on the shape of the angular distribution before and after the crossing. As such, it cannot be employed accurately in the context of hydrodynamic simulations which rely on moment-based neutrino transport schemes because of their intrinsic limitations in the reconstruction of the neutrino angular distributions in post-processing (cf. Refs. [27, 78]). For initial ELN-XLN configurations with large neutrino-antineutrino asymmetry ($\langle D_{0}^{z}\rangle\gtrsim 0.25$), we find that the angular crossing is erased by flavor conversion. However, for small neutrino-antineutrino asymmetries ($\langle D_{0}^{z}\rangle\lesssim 0.20$) and more pronounced forward-peaked distributions, new crossings can be created as the “flipped” region of the angular distribution changes sign. For small neutrino-antineutrino asymmetry ($\langle D_{0}^{z}\rangle\lesssim 0.25$), flavor evolution is responsible for driving the system towards flavor equipartition. Our results suggest that the final flavor configuration strongly depends on $\langle D_{0}^{z}\rangle$. Thus, it cannot be predicted relying on arbitrary truncations of the multipole expansion of $D_{v}^{z}$. These findings overcome earlier approaches proposed in the literature, highlighting their limitations.
Figure 5: Left: Average ELN-XLN distribution at $t=900\mu^{-1}$ for the flavor configuration C4 and three different initial seed perturbations. Independent of the perturbation, the $v>0$ part of the ELN-XLN distribution changes sign as a result of fast flavor conversion. The initial perturbation mildly affects the shape of the distribution in the angular range with $v<0$. Right: Space-averaged and angle-integrated $\nu_{e}$ survival probability as a function of time for the flavor configuration C4. The initial perturbation affects the onset of flavor conversion and the early evolution of the system, but the effect of the perturbation on the asymptotic $\langle P_{ee}\rangle$ is relatively small.
@@ Acknowledgements.
We thank Meng-Ru Wu and Masamichi Zaizen for useful discussions and comments on our manuscript. This project has received support from the European Union (ERC, ANET, Project No. 101087058) and the German Research Foundation (DFG) through the Collaborative Research Center “Neutrinos and Dark Matter in Astro- and Particle Physics (NDM),” Grant No. SFB-1258-283604770. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The Tycho supercomputer hosted at the SCIENCE HPC Center at the University of Copenhagen was used to support our numerical simulations.
Appendix A Dependence of the steady-state flavor configuration on the choice of the initial perturbation
In this paper, flavor conversion is triggered by a Gaussian perturbation described in Eq. (10). This choice was motivated by its simple analytical form and its narrow Fourier spectrum. Other studies of neutrino quantum kinetics instead generate random seeds of flavor coherence, see e.g. Refs. [33, 56]. In this appendix, we investigate whether the steady-state flavor configuration depends on the choice of the initial perturbation. To this purpose, we consider the evolution of our initial configuration C4 (cf. Table 1) under three different initial conditions: two Gaussian perturbations of width $s_{p}=9\mu^{-1}$ (used in Secs. IV and V) and $s_{p}=27\mu^{-1}$, and a random perturbation $\rho_{v,xe}(0,r)=10^{-5}\epsilon(r)$, with $\epsilon(r)\in[0,1]$.
Figure 5 shows the angular distributions at $t_{f}=900\mu^{-1}$ (left panel) and the time evolution of the angle-integrated and space-averaged $\nu_{e}$ survival probability (right panel). The evolution of $\langle P_{ee}\rangle$ in the cases with Gaussian perturbations is qualitatively the same, whereas the random initial seed is responsible for a sharper transition towards the quasi-steady state. However, we find small differences in the final value of $\langle P_{ee}\rangle$. Hence, Eq. (14) remains a better prediction of the steady state than (15) and Eqs. (16). In particular, the angular range with $v>v_{c}$ is not driven to equipartition in any of our examples, as visible in the left panel of Fig. 5; rather, the initial perturbation mainly affects the shape of the angular distribution in the range $v<0$.
Reference [33] pointed out that the choice of the perturbation seed can alter the evolution of coherent flavor structures and found that point-like perturbations may result in larger flavor conversion than random perturbations (cf. their Fig. 6). Our results suggest that, although random perturbation seeds lead to slightly higher average survival probabilities at late times, equipartition on one side of the ELN-XLN crossing is not a generic outcome of fast instabilities. Rather, it is a more common outcome in ensembles that have a shallow initial ELN-XLN crossing. In general, we expect the effect of the initial perturbation on the asymptotic flavor configuration to be negligible, in agreement with findings obtained for slow flavor conversion [83, 84, 85].